Calculate the pH to the correct number of significant figures for solutions with the following concentrations of H3O+ or OH? .

Part A

Part complete

[H3O+] = 2.5×10?5 M

SubmitPrevious AnswersRequest Answer

Part B

Part complete

[OH?] = 10?3 M

pH =

SubmitPrevious AnswersRequest Answer

Part C

Part complete

[H3O+] = 3.50×10?9 M

pH =

SubmitPrevious AnswersRequest Answer

Part D

Part complete

[H3O+] = 9×10?3 M

pH =

SubmitPrevious AnswersRequest Answer

Part E

Part complete

[OH?] = 15 M

pH =
pH =
Answer #1

Answer - Part A

[H3O+] = 2.5×10?5 M

pH = -log[H3O+]

pH = -log[2.5*10-5]

pH = 4.6

Part - B

[OH?] = 10?3 M

pOH = -log[10-3]

pOH = 3

pH = 14-3 = 11

pH+pOH = 14

Part C. [H3O+] = 3.5*10-9 M

pH = -log[H3O+]

pH = -log[ 3.5*10-9]

pH = 8.4559

Part D -

[H3O+] = 9*10-3 M

pH = -log[H3O+]

pH = -log[ 9*10-3]

pH = 2.0457

Part E - [OH-] = 15M

Here it seems mistake beacause pOH beaceme (-1.176), which is not possible.

let us assume,

[OH-] = 0.15M

pOH = -log[0.15]

pOH = 0.823

pH = 14-0.823 = 13.176